How far from a -7.40 μC point charge must a 2.50 μC point charge be placed in order for the electric potential energy of the pair of charges to be -0.600 J ? (Take the energy to be zero when the charges are infinitely far apart.) Express your answer in meters
Answer #1

let

q1 = -7.4 micro C

q2 = 2.5 micro C

let d is thebrequired distance.

we know, potential energy of the pair of charges U = k*q1*q2/d

d = k*q1*q2/U

= 9*10^9*(-7.4*10^-6)*(2.5*10^-6)/(-0.6)

= 0.2775 m